OpenStudy (caitlinnr14):

Please HELP WILL GIVE MEDAL! [3.06] Given the function f(x) = −2x + 8, find x if f(x) = 22. −36 −7 52 −15 (please explain)

OpenStudy (anonymous):

It's a function. Just plug 22 in everywhere you see x in the function. 22 is called the argument.

OpenStudy (anonymous):

Be careful as they get more complicated...if the argument is negative, I like to put parentheses around it before I plug it in...this makes me think about order of operations.

OpenStudy (zpupster):

not quite, I think f(x) = 22 so replace f(x) with 22 and then solve for x

OpenStudy (caitlinnr14):

oh ok wait who is correct

OpenStudy (anonymous):

f(x) = 2(x+6) = 22 2(x+6) = 22 2x + 12 = 22 2x = 10 x = 5

OpenStudy (zpupster):

well some people like to think f(x) as y so y = -2x + 8 and y = 22

OpenStudy (anonymous):

Oh whoops! Sorry...xpupster is right. I was reading too fast, just answered a similar problem

OpenStudy (zpupster):

there is no 5 in your multiple choice try to solve the way I showed you

OpenStudy (caitlinnr14):

ok

OpenStudy (caitlinnr14):

wait there os no 5 @kathy0514

OpenStudy (anonymous):

Follow Kathy's model but use the function in your problem. Notice that she used a different function

OpenStudy (caitlinnr14):

−7

OpenStudy (anonymous):

Sounds good :)

OpenStudy (zpupster):

good job cait

OpenStudy (anonymous):

great job!

OpenStudy (caitlinnr14):

yay thxx guys

OpenStudy (anonymous):

bye